Mess:

Asymptotes of x^2 - y^2 = 1

BobaFett:

y=+ or - (b/a(x-h)+k) y=1/1(x-0) y= -1/1(x-0) y=x y=-x

Mess:

Could you help me with a couple more?

BobaFett:

i can try

Mess:

Thanks

Mess:

Asymptotes of f(x) = √2x^2+1/3x-5

BobaFett:

vertical asymptote of √2x^2+1/3x-5 is x=5/3 horizontal asymtotes of √2x^2+1/3x-5 is y=√2/3 and y= -√2/3

Mess:

Thank you @bobafett

BobaFett:

and no problem @Mess

Mess:

Thanks again @bobafett